the ratio of william's age to harry's age is 2:5
five years ago the ratio was 1:3
in how many years time will the ratio be 4:9?

Answer by lwsshak3(11628)

the ratio of william's age to harry's age is 2:5
five years ago the ratio was 1:3
in how many years time will the ratio be 4:9?
***
at present:
w=william's age
h=harry's age
w/h=2/5
w=2h/5
x=years ratio of w/h will be 4/9
..
5 yrs ago:
w-5=william's age
h-5=harry's age
(w-5)/(h-5)=1/3
(2h/5-5)/(h-5)=1/3
h-5=3(2h/5-5)
h-5=6h/5-15
h=6h/5-10
5h=6h-50
h=50
w=2h/5=100/5=20
..
(20+x)/(50+x)=4/9
4(50+x)=9(20+x)
200+4x=180+9x
5x=20
x=4
in how many years time will the ratio be 4:9? 4